# 6DOPE-GS: Online 6D Object Pose Estimation using Gaussian Splatting

### v1

@@ 1 Introduction
Computer vision systems for our 3D world are expected to advance beyond static scenes and structured multi-camera setups to more challenging real-world applications. Precise tracking and accurate reconstruction of objects allows capturing essential spatial and structural information, essential for downstream tasks such as robotic manipulation [10, 56], augmented reality [57, 78], automation [26, 15], etc.
The majority of 6D object pose estimation and tracking methods, whether for seen or unseen objects, have primarily used model-based techniques. Several approaches [41, 81, 28, 30] use CAD models rendered from various angles during training and perform feature matching at inference time for rapid pose estimation. FoundationPose [72] leverages synthetic training data to perform state-of-the-art instance-level pose estimation using either a CAD model or a small set of reference images annotated with the object poses. Notably, there has been exciting progress in zero-shot, model-free methods over the past few years [69, 71]. BundleSDF [71] operates in a model-free manner by jointly optimizing a “neural object field” and the object poses by learning a 3D Signed Distance Field representation while concurrently running a global pose graph optimization. However, despite BundleSDF’s reported near real-time pose optimization capabilities ($\sim$10Hz), the neural object field training is far from real-time, which limits the average tracking frequency to $\sim$0.4Hz. The significant computational overhead associated with training the neural object field hinders its applicability in live dynamic scenarios, where rapid pose updates are crucial.
@@ 4.2 Metrics & Baselines
We assess pose estimation accuracy using the area under the curve percentage for the ADD and ADD-S (ADD-Symmetric) metrics [20, 74]. While these are typically computed using the object model, similarly to [69, 71], we use the ground-truth pose of the object in the first frame for calculating the metrics. For object reconstruction, we measure the Chamfer distance between the reconstructed and ground-truth object meshes in object coordinates. We compare our approach against other fully model-free RGB-D baselines, namely BundleTrack [69] and BundleSDF [71], that we ran using their open-source implementations [^1] https://github.com/NVlabs/BundleSDF. We additionally report the performance of other approaches mentioned in the leaderboard of [71]. We use the same object masks as input to all methods for a fair comparison.
@@ 4.3 Results
NICE-SLAM [80] 23.41 12.70 6.13
SDF-2-SDF [55] 28.20 14.04 2.61
DROID-SLAM [61] 46.39 34.68 4.63
MaskFusion [52] 41.88 35.07 2.34
BundleTrack [69] 93.01 87.26 2.81
BundleTrack 92.54 84.91 -
BundleSDF [71] 93.77 86.95 1.16
BundleSDF 92.82 84.28 1.23
6DOPE-GS 93.79 87.82 0.83
Table 1: Comparison on the YCBInEOAT Dataset. ADD and ADD-S are reported as AUC percentages (0 to 0.1m), and reconstruction accuracy is measured by Chamfer loss. All baseline results are the officially reported ones, except for the ones marked with an asterisk $(^{\ast})$, which we reproduced from the original implementation available on GitHub by the authors.
Method Pose Reconstruction
ADD-S (%) $\uparrow$ ADD (%) $\uparrow$ CD (cm) $\downarrow$
NICE-SLAM [80] 22.29 8.97 52.57
SDF-2-SDF [55] 35.88 16.08 9.65
KinectFusion [40] 25.81 16.54 15.49
DROID-SLAM [61] 64.64 33.36 30.84
BundleTrack [69] 92.39 66.01 52.05
BundleTrack 93.96 77.75 -
BundleSDF [71] 96.52 92.62 0.57
BundleSDF 94.86 89.56 0.58
6DOPE-GS 95.07 84.33 0.43
Table 2: Comparison on the HO3D Dataset. ADD and ADD-S are reported as AUC percentages (0 to 0.1m), and reconstruction accuracy is measured by Chamfer loss. All baseline results are the officially reported ones, except for the ones marked with an asterisk $(^{\ast})$, which we reproduced from the original implementation available on GitHub by the authors.
As seen in Tables 2 and 2, we outperform previous SLAM-based approaches as well as BundleTrack [69]. Since the sequences in YCBInEOAT (Table  2) have smoother motions with less diverse viewpoints, most approaches fare equally well as we do not have large occlusions or jumps which can cause erroneous coarse pose initialization. However, we achieve better reconstruction at a sub-centimeter level compared to other approaches on the YCBInEOAT dataset. Some examples of the pose estimation and reconstruction are shown in Fig. 2. It can be seen that 6DOPE-GS achieves accurate pose tracking and accurately reconstructs the surface appearance and geometry as more of the object becomes visible over time.
The HO3D dataset is more challenging, as there are more occlusions from human hands as well as a large range of rotational motions which cause ambiguities in the coarse pose initialization. Although we perform better than BundleSDF [71] in terms of the ADD-symmetric score, we do not compare favorably in terms of the absolute score.
This shortcoming mainly stems from the different rendering pipelines, and subsequently the gradient updates for the pose estimation. The ray-tracing-based approach of learning the neural object field via an SDF in BundleSDF enables sampling the rays in an i.i.d. manner that is more conducive for optimization. The differentiable rendering in Gaussian Splatting, on the other hand, requires the entire image to be rendered and subsequently optimized for each keyframe pose individually. While in theory, this limitation can be improved with a mini-batch-styled processing, doing so, would increase both the computational time and memory usage, potentially causing issues in the gradient accumulation due to the large number of Gaussian particles.
Figure 3: Comparison between speed and performance for different approaches on the HO3D dataset. While BundleSDF achieves high performance, it comes at the cost of speed. On the other hand, 6DOPE-GS achieves a favorable tradeoff between speed and performance. $(^{\ast})$ implies that results are based on the implementation available on GitHub by the authors.
4.4 Temporal Efficiency
To evaluate the temporal efficiency of different approaches, we compare the tradeoff between the performance and the average processing time per frame for the different approaches on the HO3D dataset. We test the approaches on a desktop with a 12th Gen Intel(R) Core(TM) i9-12900KF CPU, 64GB RAM, equipped with an NVIDIA GeForce RTX 4090 GPU. We explore the performance of two more versions of BundleSDF [71] that reduce the processing time. In the first version, named “BundleSDF-async”, we deactivate the synchronization between the threads performing neural object field learning and the online pose graph optimization. Although this approach introduces potential variability in pose accuracy, it reduces latency and achieves more real-time performance. In the second version, which we call “BundleSDF-lite”, we reduce the number of optimization steps while learning the neural object field, enabling faster synchronization between the threads.
From Fig. 3, we observe that the high pose tracking accuracy of BundleSDF [71] comes at a high computational cost. Since the pose tracking thread waits to get synchronized with the neural object field thread, it requires an average processing time of 2.1 seconds. One surprising result is the fact that BundleSDF-async achieves more accurate performance than BundleSDF-lite even though BundleSDF-async runs the pose estimation without waiting for the neural object field. This result exhibits the dependence of the pose graph optimization on having accurate keyframe poses. While the neural object field training in BundleSDF-async eventually yields more accurate poses (although at a delayed timestep) than BundleSDF-lite, the latter is unable to provide accurate pose estimates given the premature termination of the optimization to achieve faster speeds. In contrast, 6DOPE-GS provides a more balanced tradeoff between speed and accuracy. We achieve competitive performance without having to compromise on speed ($\sim$5$\times$ speedup compared to BundleSDF) as a result of the rapid convergence of the Gaussian Object Field optimization.
4.5 Ablations
We assessed our design choices on the HO3D dataset, chosen for its variety of scenarios, the results of which are shown in Table 3. Performance was reduced without dynamic keyframe selection (Ours w/o KF selection) due to the retention of inaccurate pose estimates during training, which introduces residual errors in the reconstruction loss and hinders pose optimization. Applying the vanilla adaptive density control (Ours w/o Pruning), where all Gaussians below a predefined threshold are removed, causes abrupt changes in the number of Gaussians. This results in significant rendering fluctuations, slowing the convergence of training. We find that our approach with the proposed additions, namely Dynamic Keyframe Selection and the Opacity Percentile-based Adaptive Density Control performs the best among all.
A visual comparison of the different ablations in reconstructing a water pitcher from HO3D can be seen in Fig. 4. This example is particularly difficult as there are large rotational motions and large occlusions. In this example, we find that neither of the ablations can accurately reconstruct the pitcher, especially the handle. Rather, they end up with two partially reconstructed handles as a result of inaccurate keyframe estimates which cause points to be added from a new keyframe with the handle visible. However, by combining the dynamic keyframe selection and the opacity percentile-based adaptive density control, our final approach reconstructs the handle more accurately.
Method Pose Reconstruction
ADD-S (%) $\uparrow$ ADD (%) $\uparrow$ CD (cm) $\downarrow$
Ours (basic) 93.52 80.25 0.44
Ours w/o KF Selection 94.44 82.40 0.41
Ours w/o Pruning 92.48 80.87 0.42
Ours (final) 95.07 84.33 0.43
Table 3: Ablation study of 6DOPE-GS on the HO3D dataset. “Ours w/o KF Selection” removes the dynamic keyframe selection strategy (Sec. 3.3) and performs joint optimization using all keyframes. “Ours w/o Pruning” replaces the Opacity Percentile-based Adaptive Density Control 3.4 and employs vanilla adaptive density control [25]. “Ours (basic)” is a basic version of 6DOPE-GS that naively uses all the keyframes and uses the vanilla adaptive density control.
Figure 4: Object reconstruction example using the different ablations over our approach.

### v2

@@ 1 Introduction
Precise tracking and accurate reconstruction of objects allows capturing essential spatial and structural information, essential for downstream tasks such as robotic manipulation [9, 56], augmented reality [57, 77], automation [26, 14], etc. The majority of 6D object pose estimation and tracking methods, whether for seen or unseen objects, have primarily used model-based techniques. Several approaches [41, 79, 28, 30, 72] use CAD models rendered from various angles during training and perform feature matching at inference time for rapid pose estimation. Such model-based approaches augmented with synthetic training data [72] have shown state-of-the-art performance instance-level pose estimation. However, doing so requires either a CAD model or a small set of reference images annotated with the object poses, which becomes tedious as the number of unseen objects increases.
On the other hand, there has been exciting progress in zero-shot, model-free methods over the past few years [69, 71] which require no additional prior information, other than an object mask. BundleSDF [71] operates in a model-free manner by jointly optimizing a “neural object field” and the object poses by learning a 3D Signed Distance Field representation while concurrently running a global pose graph optimization. However, despite reporting near real-time pose optimization capabilities ($\sim$10Hz), the neural object field training is far from real-time [^1] https://github.com/NVlabs/BundleSDF/issues/122, which limits the average tracking frequency to $\sim$0.4Hz. The significant computational overhead associated with training the neural object field hinders its applicability in live dynamic scenarios, where rapid pose updates are crucial.
@@ 4.2 Metrics & Baselines
We evaluate the performance of different methods based on three key metrics: 6-DoF object pose tracking accuracy, 3D reconstruction accuracy, and computational efficiency. Pose estimation accuracy is assessed using the area under the curve percentage for the ADD and ADD-S (ADD-Symmetric) metrics [20, 74]. For object reconstruction, we measure the Chamfer distance between the reconstructed and ground-truth object meshes in object coordinates. Computational efficiency is evaluated based on the average processing time per frame.
We compare several SLAM-based approaches, including DROID-SLAM (RGBD) [61], NICE-SLAM [78], KinectFusion [40], and SDF-2-SDF [55]. We report the performance of other approaches listed on the leaderboard of [71]. We also evaluate the recent 3DGS-SLAM-based method MonoGS [35], along with BundleTrack [69] and BundleSDF [71], using their open-source implementations [^2] https://github.com/NVlabs/BundleSDF, [^3] https://github.com/muskie82/MonoGS with optimized parameters. Each method takes as input an RGBD video and a per-frame mask indicating the object of interest. For a fair comparison, we evaluate MonoGS using only RGBD input.
Method Pose Reconstruction Efficiency
ADD-S (%) $\uparrow$ ADD (%) $\uparrow$ CD (cm) $\downarrow$ ATPF(s) $\downarrow$
NICE-SLAM [78] 23.41 12.70 6.13 n.a.
SDF-2-SDF [55] 28.20 14.04 2.61 n.a.
DROID-SLAM [61] 46.39 34.68 4.63 n.a.
MaskFusion [52] 41.88 35.07 2.34 n.a.
MonoGS(RGB-D) [35] 20.16 15.32 2.43 0.29
BundleTrack [69] 92.54 84.91 - 0.21
BundleSDF [71] 92.82 84.28 0.53 0.82
6DOPE-GS 93.79 87.82 0.15 0.22
Table 1: Comparison on the YCBInEOAT Dataset. ADD and ADD-S are reported as AUC percentages (0 to 0.1m), and reconstruction accuracy is measured by Chamfer loss. ATPF is the average processing time per frame (n.a. indicates unavailable data).
Method Pose Reconstruction Efficiency
ADD-S (%) $\uparrow$ ADD (%) $\uparrow$ CD (cm) $\downarrow$ ATPF(s) $\downarrow$
NICE-SLAM [78] 22.29 8.97 52.57 n.a.
SDF-2-SDF [55] 35.88 16.08 9.65 n.a.
KinectFusion [40] 25.81 16.54 15.49 n.a.
DROID-SLAM [61] 64.64 33.36 30.84 n.a.
MonoGS(RGB-D) [35] 2.81 1.82 22.09 0.36
BundleTrack [69] 93.96 77.75 - 0.29
BundleSDF [71] 94.86 89.56 0.58 2.10
6DOPE-GS 95.07 84.33 0.41 0.24
Table 2: Comparison on the HO3D Dataset. ADD and ADD-S are reported as AUC percentages (0 to 0.1m), and reconstruction accuracy is measured by Chamfer loss. ATPF is the average processing time per frame (n.a. indicates unavailable data).
@@ 4.3 Results
As seen in Tables 2 and 2, we outperform previous SLAM-based approaches as well as BundleTrack [69]. In the YCBInEOAT dataset (Table 2), where object motions are relatively smooth with less viewpoint diversity, most approaches perform similarly due to the absence of large occlusions or abrupt motion discontinuities that could lead to erroneous coarse pose initialization. We also compare our method to MonoGS [35], a state-of-the-art SLAM approach that employs Gaussian splitting for camera pose relocalization. On the YCBInEOAT dataset, MonoGS performs suboptimally due to limited geometric and textural supervision, leading to degraded object pose refinement, particularly in low-texture regions. On the HO3D dataset (Table 2), which presents more challenging scenarios with complex hand-object interactions and rapid motion variations, all baselines struggle with accurate tracking due to accumulating errors. In contrast, our pose graph optimization and keyframe selection enhance robustness in the coarse pose initialization and pose tracking efficiency, which further results in superior reconstruction at a sub-centimeter level. 6DOPE-GS maintains a strong balance between pose accuracy and temporal efficiency, making it well-suited for real-time applications.
While 6DOPE-GS performs better than BundleSDF [71] in the ADD-symmetric score, we still fall behind in the absolute score on the more challenging HO3D dataset. One possible reason is the occlusions in HO3D, which reduce the supervisory signals needed to optimize Gaussian Particles and refine pose estimates. In contrast, BundleSDF uses mini-batch SDF rendering, allowing multiple correlated optimization steps and achieving better performance. However, the high computational cost of BundleSDF limits its applicability for real-time scenarios. Our method, by contrast, provides a well-balanced trade-off between efficiency and accuracy, making it a more practical choice for real-world applications. Some examples of the precise pose tracking and accurate surface and appearance reconstruction of 6DOPE-GS are shown in Fig. 2.
Figure 3: Comparison between temporal efficiency and performance for different approaches on the HO3D dataset. While BundleSDF achieves high performance, it comes at the cost of speed. On the other hand, 6DOPE-GS achieves a favorable tradeoff between speed and performance.
4.4 Temporal Efficiency
To evaluate the temporal efficiency of different approaches, we compare the tradeoff between the performance and the average processing time per frame for the different approaches on the HO3D dataset. We test the approaches on a desktop with a 12th Gen Intel(R) Core(TM) i9-12900KF CPU, 64GB RAM, equipped with an NVIDIA GeForce RTX 4090 GPU. We explore the performance of two more versions of BundleSDF [71] that reduce the processing time. In the first version, named “BundleSDF-async”, we deactivate the synchronization between the threads performing neural object field learning and the online pose graph optimization. While this introduces potential variability in pose accuracy, it reduces latency and achieves more real-time performance. In another version, “BundleSDF-lite”, we reduce the number of optimization steps for learning the neural object field, enabling faster synchronization between the threads.
From Fig. 3, we observe that the high pose tracking accuracy of BundleSDF [71] comes at a high computational cost. Since the pose tracking thread waits for the neural object field thread to converge and subsequently synchronize the object pose and reconstruction estimates, it requires an average processing time of 2.1 seconds. Surprisingly, BundleSDF-async (without the sync between the threads) achieves better performance than BundleSDF-lite even though BundleSDF-async runs the pose estimation without waiting for the neural object field. This highlights the dependence of the pose graph optimization on accurate keyframe poses. While the neural object field training in BundleSDF-async yields more accurate poses (although at a delayed timestep) than BundleSDF-lite, the pose estimation of the latter diverges given the premature termination of the optimization to achieve faster speeds. In contrast, 6DOPE-GS provides a balanced tradeoff between speed and accuracy. We achieve competitive performance without having to compromise on speed ($\sim$5$\times$ speedup over BundleSDF) as a result of the rapid convergence of the Gaussian Object Field optimization.
4.5 Ablation Study
We assessed our design choices on both the HO3D and YCBInEoat dataset, chosen for its variety of scenarios, the results of which are shown in Table 3. Ours (basic) is a simplified version of 6DOPE-GS that naively uses all keyframes and employs vanilla adaptive density control. Ours w/o KF Selection removes the dynamic keyframe selection strategy (Sec. 3.3) and performs joint optimization using all keyframes. Ours w/o Pruning replaces the Opacity Percentile-based Adaptive Density Control 3.4 and employs vanilla adaptive density control [25]. We also compare 2DGS and 3DGS representation, Ours (3DGS) replaces the 2D Gaussian representation with a 3D Gaussian representation for pose estimation and reconstruction.
Performance was reduced without dynamic keyframe selection (Ours w/o KF selection) due to the retention of inaccurate pose estimates during training, which introduces residual errors in the reconstruction loss and hinders pose optimization. Applying the vanilla adaptive density control (Ours w/o Pruning), where all Gaussians below a predefined threshold are removed, causes abrupt changes in the number of Gaussians. This results in significant rendering fluctuations, slowing the convergence of training. The pose accuracy and reconstruction quality of 3DGS (Ours (3DGS)) are inferior to 2DGS. This can be attributed to the lack of regularization on the normal and depth in 3DGS, causing the Gaussians to deviate from the object surface and consequently degrading the reconstruction quality. We find that our approach with the proposed additions, namely Dynamic Keyframe Selection and the Opacity Percentile-based Adaptive Density Control performs the best among all.
@@ 4.5 Ablation Study
HO3D Ours (basic) 93.52 80.25 0.44
Ours w/o KF Selection 94.44 82.40 0.42
Ours w/o Pruning 92.48 80.87 0.44
Ours (3DGS) 92.51 79.49 0.47
Ours (final) 95.07 84.33 0.41
YCBInEOAT Ours (basic) 92.74 85.15 0.22
Ours w/o KF Selection 93.03 86.40 0.19
Ours w/o Pruning 92.64 86.22 0.20
Ours (3DGS) 91.18 85.29 0.41
Ours (final) 93.79 87.82 0.15
Table 3: Ablation study of critical design choices
4.6 Realtime Results